A 12.0 L balloon at 84 C is cooled until it becomes 5.0 L. To what temperature was the balloon cooled?


1
Expert's answer
2022-05-27T14:57:04-0400

Given:

V1 = 12.0 L

T1 = 84°C + 273.15 = 357.15 K

P = constant

V2 = 5.0 L

T2 = unknown


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting temperature, solve the equation for T2:

T2 = V2T1 / V1

T2 = (5.0 L × 357.15 K) / (12.0 L) = 148.81 K

T2 = 148.81 K (or −124.34°C)


Answer: The balloon was cooled to −124.34°C